OpenStudy (anonymous):

A boy drops a ball from an observation tower. The ball hits the ground in 5.0 s. What is the ball's velocity at the time of impact? A. 0 m/s B. 98 m/s C. 49 m/s D. 123 m/s **not quite sure how to solve this :( thanks!!!

OpenStudy (abhisar):

In case of dropping the initial velocity is always 0m/s acceleration due to gravity is g = 9.8m/s^2 Use the equation \(\sf v=u+gt\) to find the value of v

OpenStudy (abhisar):

u=initial velocity

OpenStudy (anonymous):

okie:) so is initial velocity 0? :/

OpenStudy (abhisar):

yes

OpenStudy (jhannybean):

Another rendition of that formula is \[v_f = v_i + a_gt\]

OpenStudy (anonymous):

so v=u+gt v=0+(9.8)(5) ?

OpenStudy (anonymous):

v=49 m/s ?

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (abhisar):

There are three equations of motion

OpenStudy (anonymous):

yay!! thanks so much :)

OpenStudy (abhisar):

\(\bullet\) v=u+at \(\bullet\) \(\sf s=ut+\frac{1}{2}at^2\) \(\bullet\) \(\sf v^2=u^2+2as\)

OpenStudy (anonymous):

okie :) i will try to remember them all haha when do i know when to apply a specific equation?

OpenStudy (abhisar):

v=final velocity u=initial velocity t=time a=acceleration

OpenStudy (abhisar):

s=distance

OpenStudy (anonymous):

okie:)

OpenStudy (anonymous):

thanks!! so i just figure out which one to use based on the info they give me?

OpenStudy (abhisar):

Yes

OpenStudy (anonymous):

okie yay!! thanks!! :D

OpenStudy (abhisar):

In projectile motion there are two components of initial velocity.... The vertical component is 0 but the horizontal component is not zero

OpenStudy (anonymous):

Ahh okay! I shall try to keep that in mind!!:)